A long straight wire carries the same current you found in part a. At what distance from the...

The magnetic field at the center of a 0.80-cm-diameter loop is 2.0mT. A long straight wire carries the same current you found in part a. At what distance from the wireis the magnetic field 2.0mT? I=13A
Answer #1

B = μi/(2πr)

0.002 = 2e-7*13/r

>>>> r = 2e-7*13/0.002 = 1.3 * 10^-3 m

answered by: MyCah